Given that VS = 3V, VBE = 0.3V, IC = 9mA, and R1 = 1kΩ, calculate the base current of the transistor Q1 using KVL. Also find the current gain ( ) of the transistor Q1


1
Expert's answer
2021-08-19T14:23:18-0400


Using KVL

"3=0.3+1000I_E\\
2.7=1000I_E\\
I_E=2.7mA\\
I_E=I_C+I_B\\
I_B=I_E-I_C=2.7-9=-6.3mA\\"

Base current is 6.3mA

Current gain beta "\beta" is given by

"\beta=\dfrac{I_C}{I_B}=\dfrac{9}{6.3}=1.4"